OpenStudy (anonymous):

Please help! Rewrite log5 (2/3) + log5 (9/10) as a single logarithm.

OpenStudy (cwtan):

if loga b+loga c it can be simplify as loga bc log5 (2/3)(9/10) log5 6/10

OpenStudy (anonymous):

log5 (2*9\3*10) log5 (3\5)

OpenStudy (anonymous):

Thanks guys. Wish I could give more than one medal :/

OpenStudy (mertsj):

\[\log_{5}\frac{2}{3}(\frac{9}{10})=\log_{5}\frac{3}{5} \]

OpenStudy (anonymous):

Good work mersj!